A rigid container is at a temperature of 112oC. When heated to 224oC, the pressure was 288 kPa. What was the initial pressure?
1
Expert's answer
2020-05-16T13:42:01-0400

nRT=PV

law ideal gas

n const of moles gas, R const., V- volume const.

initial To Po

end T1 P1

T1/T0=P1/P0 (one of three gas's law (Lussak/Bouler))

T1=224 * C=497 K

T0=112 * C=385 K

P1=288 kPa

P0=223,09 kPa